ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

6. Topology

1991 MS Classification: 55, 19 Casacuberta, Carles, Universitat Autònoma de Barcelona, Spain A universal space for plus-constructions We exhibit a two-dimensional, acyclic, Eilenberg–Mac Lane space W such that, for every space X, the plus-construction X^{+} with respect to the largest perfect subgroup of π_{1}(X) coincides, up to homotopy, with the W - nullification of X; that is, the natural map X → X^{+} is homotopy initial among maps X → Y where the based mapping space from W to Y is weakly contractible. This is joint work with A. J. Berrick. Reference: A. J. Berrick and C. Casacuberta, A universal space for plusconstructions, Topology, to appear in 1998.